A study found that the mean waiting time to see a physician at an outpatient clinic was 40 minutes
with a standard deviation of 28 minutes. Use Excel to find each probability. (a) What is the probability of more than an hour’s wait? (b) Less than 20 minutes? (c) At least 10 minutes.

Answer by stanbon(75887)

A study found that the mean waiting time to see a physician at an outpatient clinic was 40 minutes
with a standard deviation of 28 minutes. Use Excel to find each probability. (a) What is the probability of more than an hour’s wait?
P(x > 60 min) = 1 - P(0<= x <= 60) = 0.3141
-------------------------------------------------
(b) Less than 20 minutes?
P(0 <= x <= 20) = 0.1610
-------------------------------------------------
(c) At least 10 minutes.
P(x >= 10) = 1- P(0<= x < 10) = 0.9346
=================================================
Cheers,
Stan H.